OpenStudy (anonymous):

the line tangent of f(x) = 2X\(2X+1) is Y = 2X +1 at point (a,b) Find b

OpenStudy (anonymous):

graph the function and the tangent line and see what the y-value of the tangent point is

OpenStudy (anonymous):

The slope of the given tangent line is 2. Find out when \(f'(x)=2\): \[f(x)=\frac{2x}{2x+1}\\ f'(x)=\frac{2}{(2x+1)^2}\] Solve for \(x\) in the following equation: \[\frac{2}{(2x+1)^2}=2\] This \(x\) tells you when the slope of the tangent line to \(f\) is 2.